OpenStudy (anonymous):

Without drawing the graph of the equation, determine how many points the parabola has in common with the x-axis and whether its vertex lies above, on, or below the x-axis. y = 4x2 – 12x + 12

OpenStudy (anonymous):

A.2 points in common; vertex above x-axis B.no points in common; vertex above x-axis C.1 point in common; vertex on x-axis D.2 points in common; vertex below x-axis

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Loser66

jimthompson5910 (jim_thompson5910):

"determine how many points the parabola has in common with the x-axis" basically, they want the x-intercepts for this bit

OpenStudy (anonymous):

So I should use slope intercept form?

jimthompson5910 (jim_thompson5910):

this is a quadratic, not a linear equation

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

^2?

jimthompson5910 (jim_thompson5910):

you use the quadratic formula to solve for x in 4x^2 – 12x + 12 = 0

OpenStudy (anonymous):

set each factor equal to zero, and then solve each factor?

OpenStudy (anonymous):

In that case is this right? 4x^2+12=0 -12x+12=0

jimthompson5910 (jim_thompson5910):

I'm not sure why you broke that up into two equations

OpenStudy (anonymous):

How else should I do it?

jimthompson5910 (jim_thompson5910):

use the quadratic formula

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

jimthompson5910 (jim_thompson5910):

in this case, a = 4, b = -12 c = 12

OpenStudy (anonymous):

... Wow that looks scary

jimthompson5910 (jim_thompson5910):

hopefully you have seen it before

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

x=12+-Square root -12^2-4*4*12/8

OpenStudy (anonymous):

\[x=12+-\sqrt{-144-192/8}\]

OpenStudy (anonymous):

Right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is b^2-4ac equal to?

OpenStudy (anonymous):

This? x=12+−square root 144−192/8

jimthompson5910 (jim_thompson5910):

no just b^2-4ac

jimthompson5910 (jim_thompson5910):

evaluate b^2-4ac when a = 4, b = -12 c = 12

OpenStudy (anonymous):

I tried! I got x=12+-Square root -12^2-4*4*12/8 :(

jimthompson5910 (jim_thompson5910):

b^2-4ac (-12)^2-4(4)(12) 144 - 192 -48 that's what I was asking for

OpenStudy (anonymous):

So was I partially correct?

jimthompson5910 (jim_thompson5910):

so we would then have \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-(-12) \pm \sqrt{(-12)^2-4(4)(12)}}{2(4)}\] \[\Large x=\frac{12 \pm \sqrt{-48}}{8}\]

OpenStudy (anonymous):

I got -12^2-4*4*12 originally

OpenStudy (anonymous):

How will that help me find the answer though @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that discriminant is negative what does that tell you?

OpenStudy (anonymous):

D. Below?

jimthompson5910 (jim_thompson5910):

what does the discriminant determine in general?

OpenStudy (anonymous):

I don't know..

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.mathwarehouse.com/quadratic/discriminant-in-quadratic-equation.php

OpenStudy (anonymous):

Oh! It tells If the solution is a real number or an imaginary number, If the solution is rational or if it is irrational, and If the solution is 1 unique number or two different numbers

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the discriminant of -48 tells us what?

OpenStudy (anonymous):

That it is irrational and there is No Real Solutions

OpenStudy (anonymous):

and theres Two Imaginary Solutions

jimthompson5910 (jim_thompson5910):

well if you have irrational solutions, then you have a solution so you can't have "no real solutions" and have them be irrational at the same time

jimthompson5910 (jim_thompson5910):

but you saying " theres Two Imaginary Solutions " is correct

jimthompson5910 (jim_thompson5910):

what does this mean for the x-intercepts?

OpenStudy (anonymous):

They will be negative? or below?

jimthompson5910 (jim_thompson5910):

a solution to an equation like 4x^2 – 12x + 12 = 0 visually corresponds to the x-intercept this is assuming you get a real numbered solution

jimthompson5910 (jim_thompson5910):

but the solutions are imaginary numbers (not real numbers)

OpenStudy (anonymous):

2 points in common?

OpenStudy (anonymous):

but I didn't get a real numbered solution :c

jimthompson5910 (jim_thompson5910):

so are there 2 x-intercepts?

OpenStudy (anonymous):

Yes, that is what I think

jimthompson5910 (jim_thompson5910):

but the discriminant is negative

jimthompson5910 (jim_thompson5910):

negative discriminant ----> no real solutions ----> ???

OpenStudy (anonymous):

No points in common?

jimthompson5910 (jim_thompson5910):

don't worry about the answer choices right now

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

if we have no real solutions, then how many x-intercepts are there?

jimthompson5910 (jim_thompson5910):

use that page I posted

OpenStudy (anonymous):

So it means 2 imaginary solutions.. thats why I thought 2 x intercepts

OpenStudy (anonymous):

I am using it still, Thanks for the page.

jimthompson5910 (jim_thompson5910):

you'd have 2 x-intercepts if you had 2 real solutions

jimthompson5910 (jim_thompson5910):

another way to think of it as a general rule the number of unique real solutions tells you how many unique x-intercepts you have

OpenStudy (anonymous):

So 2 imaginary is no x intercepts?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

there are no x-intercepts

jimthompson5910 (jim_thompson5910):

so the graph doesn't cross the x axis

OpenStudy (anonymous):

but.. all my choices have something to do with the vertex on X

OpenStudy (anonymous):

Above, Below, or On... But you said it doesn't cross

jimthompson5910 (jim_thompson5910):

I'm not talking about the vertex at the moment

jimthompson5910 (jim_thompson5910):

I'm just referring to the x-intercepts

OpenStudy (anonymous):

Oh, umm so the descriminate is -48 and there is no X intercepts

jimthompson5910 (jim_thompson5910):

remember they said "determine how many points the parabola has in common with the x-axis" so they're saying "find the number of x-intercepts"

OpenStudy (anonymous):

But there is none

OpenStudy (anonymous):

I am sorry... but I don't get how this will help me get points or a vertex :(

OpenStudy (anonymous):

Intercept has nothing to do with vertex right?

jimthompson5910 (jim_thompson5910):

so having no x-intercepts is their way of saying there are no points that are in common with the x axis

OpenStudy (anonymous):

Ok, and the vertex is above the X axis?

jimthompson5910 (jim_thompson5910):

based on the answer choices, it has to be but it's always a good idea to know how to find the vertex

jimthompson5910 (jim_thompson5910):

this page gives a good example http://www.wikihow.com/Find-the-Vertex-of-a-Quadratic-Equation

OpenStudy (anonymous):

I posted this question an hour ago Thanks for helping me through it c:

jimthompson5910 (jim_thompson5910):

you're welcome